I would need the steps A 3000 lb car travels a quarter of a circle of diameter 50m within 4 seconds at a constant speed. If the car were to travel just a little faster, it would skid. What is the coefficient of static friction between the car and the road
Answer #1

Here,

speed of car = circumference/time

speed of car = pi * 50/(4 * 4)

speed of car , v = 9.83 m/s

radius , r = 50/2 = 25 m

let the coefficient of static friction is u

balancing force

u *m * g = m * v^2/r

9.8 * u = 9.83^2/(25)

u = 0.39

the coefficient of static friction is 0.39